Graffiti Advertising, Inc., reported the following financial statements for the last two years.
  

2016 Income Statement
Sales $ 567,200
Costs of goods sold 274,005
Selling and administrative 124,729
Depreciation 54,572
EBIT $ 113,894
Interest 19,384
EBT $ 94,510
Taxes 37,804
Net income $ 56,706
Dividends $ 10,000
Addition to retained earnings $ 46,706
GRAFFITI ADVERTISING, INC.
Balance Sheet as of December 31, 2015
Cash $ 13,360 Accounts payable $ 9,500
Accounts receivable 18,990 Notes payable 14,504
Inventory 13,798 Current liabilities $ 24,004
Current assets $ 46,148
Long-term debt $ 136,480
Net fixed assets $ 344,546 Owner's equity $ 230,210
Total assets $ 390,694 Total liabilities and owners’ equity $ 390,694
GRAFFITI ADVERTISING, INC.
Balance Sheet as of December 31, 2016
Cash $ 14,346 Accounts payable $ 10,516
Accounts receivable 21,095 Notes payable 16,470
Inventory 22,758 Current liabilities $ 26,986
Current assets $ 58,199
Long-term debt $ 152,400
Net fixed assets $ 406,307 Owner's equity $ 285,120
Total assets $ 464,506 Total liabilities and owners’ equity $ 464,506

a. Calculate the operating cash flow. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Operating cash flow            $
  
b. Calculate the change in net working capital. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Change in net working capital            $
  
c. Calculate the net capital spending. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Net capital spending            $
  
d. Calculate the cash flow from assets. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow from assets            $

e. Calculate the cash flow to creditors. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to creditors            $

f. Calculate the cash flow to stockholders. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
  
Cash flow to stockholders            $
Answer #1
Graffiti Advertising Inc.
Operating Cash Flow: Earnings before interest and tax+depreciation-Income Tax
a) Earnings before interest and tax=(A) $                          1,13,894.00
Depreciation=(B) $                             54,572.00
Income Tax=(C ) $                             37,804.00
Operating Cash Flow=(A)+(B)-(C ) $                          1,30,662.00
b) Change in Working Capital=Ending Working Capital-Beginning Working Capital
Working Capital at the end=(Ending Current Assets-Ending Current Liabilities)
Working Capital at the beginning=(Beginning Current Assets-Beginning Current Liabilities)
Ending Current Assets $                             58,199.00
Ending Current Liabilities $                             26,986.00
Working Capital at the end=($58199-$26986)=(A) $                             31,213.00
Beginning Current Assets $                             46,148.00
Beginning Current Liabilities $                             24,004.00
Working Capital at the beginning=($46148-24004)=(B) $                             22,144.00
Change in Net working capital=($31213-$22144)=(A)-(B) $                                9,069.00
c ) Capital Spending=Ending Net Fixed Assets-Beginning Net Fixed Assets+Depreciation
Ending Net Fixed Assets=(A) $                          4,06,307.00
Beginning Net Fixed Assets=(B) $                          3,44,546.00
Depreciation=(C ) $                             54,572.00
Capital Spending=(A)-(B)+(C ) $                          1,16,333.00
d ) Cash Flow from Assets=OCF-Change in Working Capital-Net capital spending
OCF=(A) $                          1,30,662.00
Change in Working Capital=(B) $                                9,069.00
Net Capital Spending=(C ) $                          1,16,333.00
Cash Flow from Assets=(A)-(B)-(C ) $                                5,260.00
e) Cash Flow to creditors=Interest Paid-Change in Long term Debt
Interest Paid=(A) $                             19,384.00
Change in long term debt=(Ending long term debt-Beginning long term debt)=($152400-136480)=(B) $                             15,920.00
Cash Flow to creditors=(A)-(B) $                                3,464.00
f) Cash flow to stockholders=Dividend Paid-Change in Equity
Dividend Paid=(A) $                             10,000.00
Change in Equity=Ending Equity-Beginning Equity-Addition to retained earnings=($285120-$230210-$46706)=(B) $                                8,204.00
Cash flow to stockholders=(A)-(B) $                                1,796.00